A frog jumps and takes 2 seconds to land. The path of the frog’s flight is best modeled by the graph of f(x) = -x2 + 2x. What is the maximum height of the jump?
Answer by Alan3354(69443)

Same problem, different numbers.
=================================================================
Determine whether there is a maximum or minimum value for f(x) = -2x^2 + 3x + 4 and find that value
---------------
The coefficient of the x^2 term is negative --> a maximum
------------
=======
The max is the vertex of the parabola at x = -b/(2a) = -3/-4 = 3/4
---
f(3/4) = -2*(3/4)^2 + 3*(3/4) + 4 = -9/8 + 9/4 + 4 = 41/8